# Score-based Generative Diffusion Models to Synthesize Full-dose FDG Brain PET from MRI in Epilepsy Patients

### v1

@@ 1 Introduction
Epilepsy, a prevalent neurological disorder characterized by recurrent seizures, necessitates precise diagnostic approaches to ensure effective management. 18fluoro-2-deoxyglucose (18F-FDG) Positron Emission Tomography (PET) imaging of brain glucose metabolism is a well-established and widely available technique for localizing epileptogenic foci within the brain. PET provides functional information by detecting metabolic changes in the brain, which are crucial for localizing epileptogenic zones (EZs), especially in cases where MRI findings are inconclusive. Meanwhile, Magnetic Resonance Imaging (MRI) is an important diagnostic tool, offering detailed anatomical images and enabling the identification of structural abnormalities such as mesial temporal sclerosis and focal cortical dysplasias that may underlie epileptic activity [1]. Research has shown that a multi-modal platform approach integrating the functional imaging of PET with the morphologic information from MRI in presurgical evaluation of epilepsy can greatly improve outcomes [2]. One study also suggests that diagnostic accuracy for EZ detection in focal epilepsy could be higher in FDG-PET/MRI compared with FDG-PET/CT [3].
@@ 2.1 Diffusion Models in Medical Imaging
Denoising Diffusion Probabilistic Models (DDPMs) have captured significant attention in recent years as a class of generative models that generate data by learning to reverse a diffusion process. These models operate in two stages: in the forward process, the original data is gradually corrupted with Gaussian noise over multiple timesteps, eventually becoming pure noise. In the backward process, the model learns to reverse this transformation step-by-step, recovering the original data distribution from the noise. Conditional inputs such as prior image or textual context can be passed into the network for all timesteps This framework has been actively explored in medical imaging due to its ability to capture complex data distributions. For instance, a DDPM with fast training and sampling was applied to T1w-to-T2 image translation using the Brain Tumor Segmentation (BraTS) dataset, with performance evaluated through metrics like PSNR and SSIM [11]. Additionally, DDPMs have shown promise in CT-to-MRI image translation [12] and for PET denoising [13].
Closely related to DDPMs, Score-Based Generative Models (SGMs) learn the gradient of the log probability density function of the data of interest, known as the score function. SGMs leverage stochastic differential equations (SDEs) to disturb, model, and sample from the data distribution, using sampling techniques such as Predictor-Corrector, Euler-Maruyama, and Heun’s methods [15, 16].
@@ 2.2 Conditional SGM
$\mathrm{L_{SGM}=E_{t}{\lambda_{t}E_{x_{0}}E_{x_{t}|x_{0},y}[\left\lVert s_{\theta}(x,y,t)-\nabla_{x}\log p_{0t}(x_{t}\mid x_{0},y)\right\rVert^{2}_{2}]}}$ (9)
$\mathrm{L_{SGM}}=E_{t}{\lambda_{t}E_{x_{0}}E_{x_{t}|x_{0},y}[\left\lVert s_{\theta}(x,y,t)+\frac{\epsilon}{\sigma_{t}}\right\rVert^{2}_{2}]}$ (10)
For the backward sampling process, our sampler follows a predictor and corrector style from [15] with the following algorithm. Starting with from Gaussian noise, the predictor uses the score from the network combining with equation (7) to iteratively update x. After each predictor step there is a fixed number of corrector steps that uses a normalized and scaled score to further improve the current x. The overall workflow of the reverse process is shown in Figure 1(b). At each timestep, the noisy prediction, the timestep information, and the conditional inputs are passed to the neural network, which outputs the gradient of the current prediction used for updates and corrections.
@@
9:     for $m=1\text{ to }M$ do $\triangleright$ Corrector step
@@
Another proposed architecture using SGM comes from Karras et al [16], whose work envisions a denoiser D reconciling different SDE schemes with the following formulation:
@@
Here is the noise level at a specific timestep. $F_{\theta}$ is the neural network to be trained on noise-disturbed data, $c_{skip}(\sigma)$ is a $\sigma$-dependent scalar modulating the skip connection, $c_{in}(\sigma)$ and $c_{out}(\sigma)$ modulates the input and output magnitude and $c_{noise}(\sigma)$ is a scalar served as a conditional input into the network. In our task, y is the MRI/PET low-dose condition, and x is the full-dose PET data, we can add our conditions into the equation:
@@
During the sampling process, we used the paper’s stochastic sampler, which averages the gradient from different timesteps.
@@
10:         $d_{i}^{\prime}\leftarrow(x_{i+1}-D_{\theta}(x_{i+1},t_{i+1}),y)/t_{i+1}$
@@ 2.3 Transformer-Unet (TransUnet)
TransUnet, developed in our previous work [9], takes in multi-contrast MRI inputs to predict full-dose FDG-PET and demonstrated strong performance based on metabolic metrics such as SUVR. As shown in Figure 1(a), the multi-contrast inputs are concatenated at the start of the network. There are 5 blocks with batch-normalization and convolution layers with stride of 2 so that the 256x256 inputs are downsampled to 8x8. Next, a transformer block is applied on the flattened inputs at this innermost layer with multi-head self-attention to capture global information. During the upsampling blocks, spatial attention, channel attention and upsampling are applied to the inputs and concatenated with the corresponding inputs from the downsampling blocks. Finally, an output convolution layer is applied to obtain the final synthetic PET slice.
@@
2.4 Patient Population
After receiving a waiver of consent from our IRB, we identified consecutive adult patients who underwent simultaneous FDG 3.0T brain PET/MRI (GE Signa, Waukesha, WI) examinations for clinical purposes between August 2022 and October 2024 with available list-mode data. Patients received a standard full-dose radiotracer. Ultralow-dose PET images were reconstructed using list-mode random count under-sampling. Specifically, for all coincidence events recorded in the list-mode data, we selected 1 event every 100 events for 1% PET simulation and passed the selected events into a 3D PET reconstruction. We tested several parameters and optimized the number of reconstruction updates to 28, beyond which no further improvement on image quality is observed and unnecessary noise may be added into the image.
@@ 2.4 Patient Population
Table 1: Table for training, validating and testing data cases
@@ 2.5 Implementation Detail
We investigated the effect of multi-sequence MRI and ultralow-dose PET using 4 different input combinations for each model: T1w only (no PET data), T1w + T2-FLAIR (no PET data), and 1% PET data combined with either T1w or T1w + T2-FLAIR (Figure 2b). For all models, we used 3 MRI slices to predict the middle PET slice. Larger numbers of surrounding slices were explored but showed limited improvement. Training was done on a 32GB GPU (Quadro GV100, Tesla V100-PCIE) with the largest batch size allowed by the memory.
@@ 2.6 Evaluation Metrics
For evaluation, brain segmentation was performed using Freesurfer’s Destrieux parcellation [18], and related small regions belonging to the same overarching larger region were grouped together (e.g., all sub-regions within frontal cortex were grouped into frontal cortex). We performed measurements in 10 region-of-interests (ROI): frontal cortex (FC), temporal cortex (TC), parietal cortex (PC), occipital cortex (OC), insular cortex (IC) cerebral white matter (CWM), deep gray matter including caudate, putamen, and globus pallidus (DGM), combined hippocampus and amygdala (HipAmy), cerebrospinal fluid (CSF), and cerebellum. Standardized Uptake Value Ratios (SUVR) were then calculated using the cerebellum as the reference region for the entire volume. The mean absolute difference in SUVR ($\Delta$SUVR Mean) between the acquired PET and synthetic PET is computed across the whole brain volume without background and across all test subjects, and the standard deviation of the voxel-wise difference ($\Delta$SUVR STD) across brain volume is also computed individually and averaged across all test subjects. This evaluates the accuracy and consistency of model performance on a voxel level.
Given the diagnostic importance of hemispheric metabolic differences in epilepsy [19], we calculated a clinically relevant metric termed the Congruence Index (CI)to assess the agreement of SUVR asymmetry between the synthesized and the acquired PET images. We first construct an Asymmetry Index (AI) [20, 21] for each ROI using equation (14), where the difference in SUVR from the left and right hemisphere is divided by the sum. As such, a positive AI indicates right-sided hypometabolism while a negative AI indicates left-sided hypometabolism [22].
@@
Next, we constructed a Congruence Index using formulation in (15) to obtain the proportion of ROIs matching the asymmetry present in the acquired PET. We identified 8 areas with laterality: FC, TC, OC, IC, PC, CWM, DGM, and HipAmy. We averaged the index across the 10 test subjects. A higher CI indicates better representation of the left-right asymmetries found in the reference full-dose image. The formulation for CI calculation is summarized in (15):
$\mathrm{CI_{synth}=\frac{1}{N_{subjects}}\sum_{subjects}\frac{\sum_{ROIs}[[AI_{ROI}(PET_{synth})\times AI_{ROI}(PET_{acquired})>0]]}{N_{ROIS}}}$ (15)
@@ 3 Results
3.1 Visual Quality of Synthetic PETs
For the zero-dose task using only MRI sequences as inputs, the 2 SGM models outperformed the TransUnet, as shown by the more accurate synthesis in the region highlighted by the yellow boxes in Figure 3. The TransUnet tended to generate oversmoothed images with fewer details, consistent with the observation made in [10], whereas score-based models may emphasize the edges and the contrast between regions with low and high metabolism. For the ultralow-dose task that included 1% dose PET images as input, the two SGM models exhibited reduced slice inconsistencies in coronal and sagittal views, as highlighted by the red boxes in Figure 3. The addition of 1% PET as an input image improved the quality of coronal and sagittal views by increasing the intensity consistency between adjacent slices. This is better seen in Figure 4, which provides a closer examination of the improvement in slice inconsistency when we move from zero-dose to ultralow-dose task. The ultralow-dose line profiles exhibit less jittering compared with the zero-dose profiles, reflecting smoother intensity transitions between slices that more closely resemble the acquired PET. Notably, the TransUnet shows the least slice inconsistency in the zero-dose task, but requires 1% PET input to generate sufficient anatomical detail. This limitation may stem from its reduced complexity, which could hinder performance with limited training data.
@@ 3.1 Visual Quality of Synthetic PETs
Figure 4: Slice inconsistency evaluation between zero-dose and ultralow-dose tasks using the line profile from each PET (shown by the green lines). On the bottom graphs, the red curve represents the acquired PET, the orange curve represents the zero-dose task, while the blue curve represents the ultralow-dose task. The blue curve is closer to the red curve, indicating improved slice intensity consistency in ultralow-dose task.
@@ 3.2 Quantitative Performances
When PET 1% data was included as input, there was improved performance for the asymmetry indices across all model types (Table 3). Specifically, the best performing CI increases to a high of 0.9 while the best CMAE is 0.197. TransUnet experiences the largest improvement compared to SGMs and is now capable of capturing smaller anatomical details (see Figure 3). ICC values were largely unchanged, still exceeding the 0.75 threshold indicative of good reliability, suggesting that ICC is less dependent on the presence of ultralow-dose inputs. Figure 5 presents these findings in a summary chart that allows clear visualization of the improvement of the CI and CMAE metrics for all models when the 1% PET image is included as an input.
@@
Furthermore, the performances on mean voxel-wise $\Delta$SUVR and mean $\Delta$SUVR standard deviation are shown in Table 4 and Figure 6, where the top half represents the zero-dose task and the lower half the ultralow-dose task. For the zero-dose task, SGM-KD showed both the lowest $\Delta$SUVR Mean and STD, meaning that although SGM-KD is not the best in asymmetry measures, it has a more balanced performance across the volume. For the ultralow-dose task, TransUnet has the lowest $\Delta$SUVR Mean and STD.
@@ 4 Discussion
Previously, Xiang et al. developed a CNN-based patch model for low-dose (25%) PET denoising that incorporated T1w MRI information [6]. Their approach used low-dose PET data acquired during a 3 min short scan extracted from a 12 min standard-dose tracer injection, which differs from the event under-sampling strategy we employed during list-mode reconstruction. In our prior work, we integrated transformer, generator, and discriminator components into a U-Net architecture to synthesize FDG-PET images from multi-modal inputs, including T1w, T1c, T2 FLAIR, and ASL in tumor patients [9]. While this approach demonstrated strong performance based on metabolic metrics such as SUVR, it relied on the availability of diverse contrast images, which may not always align with clinical acquisition protocols. In contrast, the current study utilizes more basic MRI inputs, improving the practicality and generalizability of the method across diverse datasets. Additionally, the availability of list-mode data provides higher flexibility to simulate different dose levels, facilitating future explorations into dose optimization.
@@
Due to the scarcity of publicly available PET datasets, particularly for epilepsy cases, we first examined prior work on PET reconstruction and synthesis in dementia studies. The availability of large ADNI datasets has enabled the development of multiple GAN-based 3D image translation approaches [24, 25, 26]. For instance, Li et al. applied patch-based 3D CNNs for MRI-to-PET translation using 19.9 million patches from ADNI [24]. Hinge et al. employed a Pix2Pix conditional GAN [25], while Sikka et al. utilized a 3D U-Net on full MR volumes to capture global context, improving Alzheimer’s classification using synthetic PET [sikka]. Similarly, Pan et al. used a cycle-consistent GAN to impute missing PET data, integrating it with MRI for AD diagnosis [27]. With a focus on clinical applications, Choi et al. trained a variational autoencoder (VAE) on normal brain PET to assist in AD and MCI diagnosis [28]. For amyloid PET reconstruction, Ouyang et al. proposed a 2.5D GAN model with task-specific perceptual loss, leveraging a pre-trained amyloid status classifier to enhance pathological feature matching [29]. Additionally, the multi-Pareto GAN introduced a multi-round generator and dynamic Pareto discriminator to iteratively denoise low-dose PET scans [30]. However, despite these advancements, the exploration of 3D diffusion architectures remains significantly underdeveloped, likely due to their high computational demands. Diffusion models hold great promise for real-world applications in image processing and acquisition. For example, a denoising diffusion null-space model was trained on high-quality PET images from a high-end scanner and integrated into a conventional PET system to enhance lower-quality scans [31]. This highlights the realistic potential of diffusion-based methods to improve PET imaging, warranting further investigation into their feasibility and impact.
@@
There are several important considerations when choosing a generative model synthesizing full dose FDG PET. TransUnet, while having the fastest sampling ($<$40s per volume), needs larger training data or more input information for zero-dose task. Thus, it is the most suitable for larger training sets containing both MRI and low-dose PET images. Of the two score-based diffusion models, SGM-KD has shorter sampling time than SGM-VP due to its use of step discretization. Specifically, we can limit the number of sampling steps to 50-100 steps ($<$37s per 8-slice inference batch, $<$400s per volume) for a satisfactory result, compared to 200-300 steps that are required for the SGM-VP ($<$220s per 8-slice inference batch, $<$2420s per volume). Further exploration on how to reduce the sampling time of SGMs are ongoing, including consistency models [33] and DPM samplers [34].

### v2

@@ 1 Introduction
Epilepsy, a prevalent neurological disorder characterized by recurrent seizures, necessitates precise diagnostic approaches to ensure effective management. 18F-fluoro-2-deoxyglucose (18F-FDG) Positron Emission Tomography (PET) imaging of brain glucose metabolism is a well-established and widely available technique for localizing epileptogenic foci within the brain. PET provides functional information by detecting metabolic changes in the brain, which are crucial for localizing epileptogenic zones (EZs), especially in cases where MRI findings are inconclusive. Meanwhile, Magnetic Resonance Imaging (MRI) is an important diagnostic tool, offering detailed anatomical images and enabling the identification of structural abnormalities such as mesial temporal sclerosis and focal cortical dysplasias that may underlie epileptic activity [1]. Research has shown that a multi-modal platform approach integrating the functional imaging of PET with the morphologic information from MRI in presurgical evaluation of epilepsy can greatly improve outcomes [2]. One study also suggests that diagnostic accuracy for EZ detection in focal epilepsy could be higher in FDG-PET/MRI compared with FDG-PET/CT [3].
@@ 2.1 Diffusion Models in Medical Imaging
DDPMs have captured significant attention in recent years as a class of generative models that generate data by learning to reverse a diffusion process. These models operate in two stages: in the forward process, the original data is gradually corrupted with Gaussian noise over multiple timesteps, eventually becoming pure noise. In the backward process, the model learns to reverse this transformation step-by-step, recovering the original data distribution from the noise. Conditional inputs such as prior image or textual context can be passed into the network for all timesteps This framework has been actively explored in medical imaging due to its ability to capture complex data distributions. For instance, a DDPM with fast training and sampling was applied to T1w-to-T2 image translation using the Brain Tumor Segmentation (BraTS) dataset, with performance evaluated through metrics like PSNR and SSIM [11]. Additionally, DDPMs have shown promise in CT-to-MRI image translation [12] and for PET denoising [13, 14]. However, based on our initial exploration, DDPMs / Latent Diffusion models did not produce satisfactory results with small datasets.
Closely related to DDPMs, Score-Based Generative Models (SGMs) learn the gradient of the log probability density function of the data of interest, known as the score function. SGMs leverage stochastic differential equations (SDEs) to disturb, model, and sample from the data distribution, using sampling techniques such as Predictor-Corrector, Euler-Maruyama, and Heun’s methods [12, 13, 14, 15, 16].
@@ 2.2 Conditional SGM
$\mathrm{Loss=E_{t}{\lambda_{t}E_{x_{0}}E_{x_{t}|x_{0},y}[\left\lVert s_{\theta}(x,y,t)-\nabla_{x}\log p_{0t}(x_{t}\mid x_{0},y)\right\rVert^{2}_{2}]}}$ (9)
$\mathrm{Loss=E_{t}{\lambda_{t}E_{x_{0}}E_{x_{t}|x_{0},y}[\left\lVert s_{\theta}(x,y,t)+\frac{\epsilon}{\sigma_{t}}\right\rVert^{2}_{2}]}}$ (10)
For the backward sampling process, our sampler follows a predictor and corrector style from [15] with the following algorithm. Starting with from Gaussian noise, the predictor uses the score from the network combining with equation (7) to iteratively update x. After each predictor step there is a fixed number of corrector steps that uses a normalized and scaled score to further improve the current x. The overall workflow of the reverse process is shown in Figure 1a. At each timestep, the noisy prediction, the timestep information, and the conditional inputs are passed to the neural network, which outputs the gradient of the current prediction used for updates and corrections.
@@
9:     for $j=1\text{ to }M$ do $\triangleright$ Corrector step
@@
Another proposed architecture using SGM comes from Karras et al. [16], which we abbreviate as SGM-KD), which envisions a denoiser D reconciling different SDE schemes with the following formulation:
@@
Here $\sigma$ is the noise level at a specific timestep. $F_{\theta}$ is the neural network to be trained on noise-disturbed data, $c_{skip}(\sigma)$ is a $\sigma$-dependent scalar modulating the skip connection, $c_{in}(\sigma)$ and $c_{out}(\sigma)$ modulates the input and output magnitude and $c_{noise}(\sigma)$ is a scalar served as a conditional input into the network. In our task, y is the MRI/PET low-dose condition, and x is the full-dose PET data, we can add our conditions into the equation:
@@
During the sampling process, we used the paper’s stochastic sampler, which averages the gradient from the current timestep and a future timestep for improved quality:
@@
10:         $d_{i}^{\prime}\leftarrow(x_{i+1}-D_{\theta}(x_{i+1},t_{i+1},y)/t_{i+1}$
@@ 2.3 Transformer-Unet (TransUnet)
TransUnet, developed in our previous work [9], takes in multi-contrast MRI inputs to predict full-dose FDG-PET and demonstrated strong performance based on metabolic metrics such as SUVR. As shown in Figure 1b, the multi-contrast inputs are concatenated at the start of the network. There are 5 blocks with batch-normalization and convolution layers with stride of 2 so that the 256x256 inputs are downsampled to 8x8. Next, a transformer block is applied on the flattened inputs at this innermost layer with multi-head self-attention to capture global information. During the upsampling blocks, spatial attention, channel attention and upsampling are applied to the inputs and concatenated with the corresponding inputs from the downsampling blocks. Finally, an output convolution layer is applied to obtain the final synthetic PET slice.
@@
2.4 Patient Population and Imaging Methods
After receiving a waiver of consent from our IRB, we identified consecutive adult patients who underwent simultaneous FDG 3.0T brain PET/MRI (GE Signa, Waukesha, WI) examinations for clinical purposes between August 2022 and October 2024 with available list-mode data. The primary clinical indication for these studies was seizure evaluation, though cases with other indications (e.g., dementia, tumor) were also present. However, only patients undergoing evaluation for epilepsy were included in the test sets, reflecting our specific focus on epilepsy in this study. All available cases, including non-epilepsy cases, were included in the training set to maximize dataset size. Dataset splits and input details are shown in Figure 2, where we identified 52 cases (9 non-epilepsy and 43 epilepsy cases). Ten and two epilepsy patients were randomly selected as test and validation cases, respectively. The remainder of the patients were used for model training.
@@ 2.4 Patient Population and Imaging Methods
Table 1: Table for training and testing data cases
@@ 2.5 Implementation Detail
We investigated the effect of multi-sequence MRI and ultralow-dose PET using 4 different input combinations for each model: T1w only (no PET data), T1w + T2-FLAIR (no PET data), and 1% PET data combined with either T1w or T1w + T2-FLAIR (Figure 2b). For all models, we used 3 MRI slices to predict the middle PET slice. Larger numbers of surrounding slices were explored but showed limited improvement. Training was done on a 32GB GPU (Quadro GV100, Tesla V100-PCIE) with the largest batch size allowed by the memory (8-16 per batch for SGMs, 64 per batch for TransUnet).
@@ 2.6 Evaluation Metrics
For evaluation, brain segmentation was performed using Freesurfer’s Destrieux parcellation [18], and related small regions belonging to the same overarching larger region were grouped together (e.g., all sub-regions within frontal cortex were grouped into frontal cortex). We performed measurements in 10 region-of-interests (ROI): frontal cortex (FC), temporal cortex (TC), parietal cortex (PC), occipital cortex (OC), insular cortex (IC) cerebral white matter (CWM), deep gray matter including caudate, putamen, and globus pallidus (DGM), combined hippocampus and amygdala (HipAmy), cerebrospinal fluid (CSF), and cerebellum. Standardized Uptake Value Ratios (SUVR) were then calculated using the cerebellum as the reference region. The mean absolute difference in SUVR ($\Delta$SUVR Mean) between the acquired PET and synthetic PET is computed across the whole brain volume across all test subjects, and the standard deviation of the voxel-wise difference ($\Delta$SUVR STD) across brain volume is also computed individually and averaged across all test subjects. This evaluates the accuracy and consistency of model performance on a voxel level.
Given the diagnostic importance of hemispheric metabolic differences in epilepsy [19], we calculated a clinically relevant metric termed the Congruence Index (CI)to assess the agreement of SUVR asymmetry between the synthesized and the acquired PET images. We first construct an Asymmetry Index (AI) [20, 21, 22] for each ROI using equation (14), where the difference in SUVR from the left and right hemisphere is divided by the sum. As such, a positive AI indicates right-sided hypometabolism while a negative AI indicates left-sided hypometabolism [22].
@@
Next, we constructed a Congruence Index using formulation in below to obtain the proportion of ROIs matching the asymmetry present in the acquired PET. We identified 8 areas with laterality: FC, TC, OC, IC, PC, CWM, DGM, and HipAmy. We averaged the index across the 10 test subjects. A higher CI indicates better representation of the left-right asymmetries found in the reference full-dose image. The formulation for CI calculation is summarized in (15):
$\mathrm{CI_{synth}=\frac{1}{N_{subjects}N_{ROIs}}\sum_{subjects,ROIs}[[AI_{ROI}(PET_{synth})\times AI_{ROI}(PET_{acquired})>0]]}$ (15)
@@
To analyze slice inconsistencies that can occur using 2D models for 3D image generation, we evaluated randomly selected slices reconstructed in the coronal orientation. From this, we computed the line profile of signal intensity along the cranial-caudad direction for the synthesized and acquired PET studies.
@@ 3 Results
3.1 Visual Quality of Synthetic Full-dose FDG-PET
For the zero-dose task using only MRI sequences as inputs, the 2 SGM models outperformed the TransUnet, as shown by the more accurate synthesis in the region highlighted by the yellow boxes in Figure 3. The TransUnet (CNN style network) tended to generate oversmoothed images with fewer details, consistent with the observation made in [12], whereas score-based models emphasize the edges and the contrast between regions with low and high metabolism. For the ultralow-dose task that included 1% dose PET images as input, the two SGM models exhibited reduced slice-to-slice inconsistencies in coronal and sagittal views, improving the quality of coronal and sagittal views (Figure 3). This is also seen in Figure 4, which provides a closer examination of the slice inconsistency improvement moving from zero-dose to ultralow-dose task. The ultralow-dose line profiles exhibit less jittering compared with the zero-dose profiles, reflecting smoother intensity transitions between slices that more closely resemble the acquired PET. Notably, the TransUnet shows the least slice inconsistency in the zero-dose task, but requires 1% PET input to generate sufficient anatomical detail. This limitation may stem from its reduced complexity, which could hinder performance with limited training data.
@@ 3.1 Visual Quality of Synthetic Full-dose FDG-PET
Figure 4: Interslice consistency evaluation between zero-dose and ultralow-dose tasks using the line profile from 1 coronal slice in each PET (intensity values taken along the green lines). On the bottom graphs, the red curve represents the acquired PET, the orange curve represents the zero-dose task, and the blue curve represents the ultralow-dose task. The blue curve is closer to the red curve, and has less high-frequency components, indicating improved slice intensity consistency for the ultralow-dose task.
Lastly, we examined the quality of the synthetic PET for a test case with left-sided temporal lobe hypometabolism based on the clinical reports. As shown in Figure 5, the left-right asymmetries are much easier to identify on the ultralow-dose synthesis as compared to the zero-dose synthesis. This is true for all three models.
Figure 5: Axial slice for one case with asymmetrically lower metabolism in the left temporal lobe based on the clinical report. We computed the mean SUVR in the left and right yellow boxes with values shown on top. We observe better agreement with the acquired PET for ultralow-dose task compared to zero-dose task.
@@ 3.2 Quantitative Performances
When PET 1% data was included as input, there was improved performance for the asymmetry indices across all model types (Table 3). Specifically, the best performing CI increases to a high of 0.9 while the best CMAE is 0.20. TransUnet experiences the largest improvement compared to SGMs and is now capable of capturing smaller anatomical details (see Figure 3). ICC values were largely unchanged, still exceeding the 0.75 threshold indicative of good reliability, suggesting that ICC is less dependent on the presence of ultralow-dose inputs. Figure 6 presents these findings in a summary chart that allows clear visualization of the improvement of the CI and CMAE metrics for all models when the 1% PET image is included as an input.
@@
Lastly, the performances on mean voxel-wise $\Delta$SUVR and mean $\Delta$SUVR standard deviation are shown in Table 4 and Figure 7, where the top half represents the zero-dose task and the lower half the ultralow-dose task. For the zero-dose task, SGM-KD showed both the lowest $\Delta$SUVR Mean and STD, meaning that although SGM-KD is not the best for asymmetry measures, it has a more balanced performance across the volume. For the ultralow-dose task, TransUnet has the lowest $\Delta$SUVR Mean and STD. We notice that although SGM-VP performs better for asymmetry measures, SGM-KD outperforms in the $\Delta$SUVR metrics. A possible explanation is the limited dataset size, which may introduce differences that would not otherwise occur. Additionally, the architecture of SGM-KD might be more effective at synthesizing pixels with average intensities, rather than those representing hypo- or hyper-metabolized regions at the distribution’s extremes. However, these performance differences may change when more data becomes available.
@@ 4 Discussion
Previously, Xiang et al. developed a CNN-based patch model for low-dose (25%) PET denoising that incorporated T1w MRI information [6]. Their approach used low-dose PET data acquired during a 3 min short scan extracted from a 12 min standard-dose tracer injection, which differs from the event under-sampling strategy we employed during list-mode reconstruction. Another prior study examining only MR-to-PET image modality synthesis integrated transformer, generator, and discriminator components into a U-Net architecture to synthesize FDG-PET images from multi-modal MR inputs, including T1w, T1c, T2 FLAIR, and ASL in brain tumor patients [9]. While this approach demonstrated strong performance based on metabolic metrics such as SUVR, it relied on the availability of diverse contrast images, which may not always align with clinical acquisition protocols. In contrast, the current study utilizes fewer and more basic MRI inputs, improving the practicality and generalizability of the method across diverse datasets. Additionally, the availability of list-mode data provides higher flexibility to simulate different dose levels, facilitating future explorations into dose optimization.
@@
Due to the scarcity of publicly available PET datasets, particularly for epilepsy cases, we first examined prior work on PET reconstruction and synthesis in dementia studies. The availability of large ADNI datasets has enabled the development of multiple GAN-based 3D image translation approaches [24, 25, 26]. For instance, Li et al. applied patch-based 3D CNNs for MRI-to-PET translation using 19.9 million patches from ADNI [24]. Hinge et al. employed a Pix2Pix conditional GAN [25], while Sikka et al. utilized a 3D U-Net on full MR volumes to capture global context, improving Alzheimer’s classification using synthetic PET [26]. Similarly, Pan et al. used a cycle-consistent GAN to impute missing PET data, integrating it with MRI for AD diagnosis [27]. With a focus on clinical applications, Choi et al. trained a variational autoencoder (VAE) on normal brain PET to assist in AD and MCI diagnosis [28]. For amyloid PET reconstruction, Ouyang et al. proposed a 2.5D GAN model with task-specific perceptual loss, leveraging a pre-trained amyloid status classifier to enhance pathological feature matching [29]. Additionally, the multi-Pareto GAN introduced a multi-round generator and dynamic Pareto discriminator to iteratively denoise low-dose PET scans [30]. Despite these advancements, the exploration of 3D diffusion architectures remains significantly underdeveloped, likely due to their high computational demands. Diffusion models hold great promise for real-world applications in image processing and acquisition. For example, a denoising diffusion null-space model was trained on high-quality PET images from a high-end scanner and integrated into a conventional PET system to enhance lower-quality scans [31]. This highlights the realistic potential of diffusion-based methods to improve PET imaging, warranting further investigation into their feasibility and impact.
@@
There are several important considerations when choosing a generative model synthesizing full dose FDG PET. TransUnet, while having the fastest sampling ($<$40s per volume), needs larger training data or more input information for zero-dose task. Thus, it is the most suitable for larger training sets containing both MRI and low-dose PET images. Of the two score-based diffusion models, SGM-KD has shorter sampling time than SGM-VP due to its use of step discretization. Specifically, we can limit the number of sampling steps to 50-150 steps ($<$37s per 8-slice inference batch,  400s per volume) for a satisfactory result, compared to 200-300 steps that are required for the SGM-VP ($<$220s per 8-slice inference batch,  2420s per volume). Further exploration on how to reduce the sampling time of SGMs are ongoing, including consistency models [33] and DPM samplers [34].